### v1

# Singular equivalence and singular locus

Abstract.
For a commutative Gorenstein Noetherian ring $R$, we construct an affine scheme $X$ solely from singularity category $D_{sg}(R)$ of $R$ such that there is a finite surjective morphism $X\rightarrow\mathrm{Spec}(R/I)$, where $\mathrm{Spec}(R/I)$ is the singular locus in $\mathrm{Spec}(R)$. As an application, for two such rings with equivalent singularity categories, we prove that the singular loci in their affine schemes have the same dimension.
1. Introduction
1.1. Background
Let $A$ be an associative algebra over a base field $k$ of characteristic zero. Its singularity category $D_{sg}(A)$ is the Verdier quotient $D^{b}(A)/\mathrm{Perf}(A)$, where $\mathrm{Perf}(A)$ is the full subcategory consisting of perfect complexes over $A$. $D_{sg}(A)$ measures the smoothness of $A$ in the sense that $A$ is homologically smooth if and only if $D_{sg}(A)$ is trivial. It was first introduced by Buchweitz in the study of algebraic representations of Gorenstein rings in [5]. Moreover, he showed that $D_{sg}(A)$ is equivalent to the Cohen-Macaulay stable category $\underline{\mathrm{CM}}(A)$ as triangulated categories when $A$ is Gorenstein. Later, in [19], Orlov rediscovered this notion from the perspective of algebraic geometry and mathematical physics, which has a deep relationship with Homological Mirror Symmetry.
In recent years, many mathematicians have studied singular categories from various perspective and made significant progress in this field, such as in tilting theory ([11, 21, 6] etc.), homological algebra ([5, 8, 15, 24, 25]), algebraic geometry ([26, 13, 3] etc.) and even in knot theory ([12]).
On the other hand, invariants under triangulated equivalence have been extensively studied and played an important role in the research of triangulated categories. It is widely known that the Hochschild cohomology and Hochschild homology are both invariants under the derived Morita equivalence. In [1], Armenta and Keller showed that the differential calculus of on the Hochschild cohomology and cohomology of an associative algebra is invariant under the derived Morita equivalence. Regarding the triangulated equivalence of singularity categories, Wang showed that the Gerstenhaber bracket structure on the Tate-Hochschild cohomology is invariant under the singular equivalence of Morita type with level (see [25]).
In noncommutative algebraic geometry, people focus more on invariants that arise from geometry. For example, Bondal and Orlov showed that for a projective variety $X$ with ample or anti-ample canonical bundle, its bounded derived category of coherent sheaves $D^{b}(X)$ recovers $X$ ([4]). Recently, in [10], Hua and Keller showed that the singularity category of a hypersurface algebra with isolated singularity recovers the algebra itself via the isomorphism between the zeroth Tate-Hochschild cohomology and the Tyurina algebra of this hypersurface.
1.2. Main result
In [20], Orlov proved that the completion of a variety along its singular locus determines its singularity category, up to the idempotent completion of a triangulated category. We want to answer the inverse question of Orlov’s result up to some extent. Unfortunately, we need some restrictive conditions on the commutative ring. Our main result is the following.
Theorem 1.1.
Let $R$ be a commutative Gorenstein Noetherian ring. Then there is an affine scheme $X$ constructed solely from the triangulated category $D_{sg}(R)$, and a finite surjective morphism $X\rightarrow\mathrm{Spec}(R/I)$, where $\mathrm{Spec}(R/I)$ is the singular locus in $\mathrm{Spec}(R)$.
In the above theorem, the singular locus is given as follows. Let $X$ be a scheme over $k$, then the singular locus of $X$ is
$\mathrm{Sing}(X):=\big{\{}\mathbf{p}\in X|X_{\mathbf{p}}\,\textup{is singular}\big{\}},$
where $X_{\mathbf{p}}$ is the localization of $X$ at point $\mathbf{p}$.
1.3. Idea of the proof
The main idea of the proof of the above result is as follows.
  1. (1)
    First, for any commutative Gorenstein Noetherian ring, we show that the coordinate ring of its singular locus is a subring of its zeroth Tate-Hochschild cohomology.
  2. (2)
    Second, we continue to show that the zeroth Tate-Hochschild cohomology is a finitely generated module over the coordinate ring of its singular locus.
  3. (3)
    At last, with some preparations (see Lemmas 4.1 and 4.2), we construct scheme $X$ and morphism $X\rightarrow\mathrm{Spec}(R/I)$, and then prove Theorem 1.1.
The paper is organized as follows. In Section 2, we introduce the notions and some properties of singularity categories, singular equivalences and DG categories. In Section 3, we prove the statements in the steps (1) and (2) above. In Section 4, we give the proof of the statement in step (3), and get the main result.
1.4. Notation and conventions
Throughout the paper, $k$ is a perfect field with characteristic zero. Unless otherwise specified, an algebra is unital over $k$, and all modules are right modules. We also assume that all commutative Noetherian rings are finitely generated over $k$ and have finite Krull dimension. For an algebra $A$, we denote by $\mathrm{mod}(A)$ the category of finitely generated $A$-modules, and by $\mathrm{proj}(A)$ its full subcategory of finitely generated projective $A$-modules. Recall that $D^{b}(A)$ is the bounded derived category of $\mathrm{mod}(A)$, and $\mathrm{Perf}(A)$ is the bounded homotopy category of complexes of $\mathrm{proj}(A)$.
2. preliminaries
2.1. Singularity category and singular equivalence
We recall the definitions of singularity category and singular equivalence for algebras.
Definition 2.1.
Let $A$ be an associative algebra. The singularity category of $A$ is defined to be the Verdier quotient $D_{sg}(A):=D^{b}(A)/\mathrm{Perf}(A)$. More precisely,
  1. (i)
    $\mathrm{Obj}\big{(}D_{sg}(A)\big{)}=\mathrm{Obj}\big{(}D^{b}(A)\big{)}$;
  2. (ii)
    for any $M^{\ast},N^{\ast}\in\mathrm{Ob}\big{(}D_{sg}(A)\big{)}$ and $f\in\mathrm{Hom}_{D^{b}(A)}(M^{\ast},N^{\ast})$ such that $\mathrm{Cone}(f)\in\mathrm{Ob}\big{(}\mathrm{Perf}(A)\big{)}$ and $f$ is invertible in $D_{sg}(A)$. Denote by $\mathcal{S}$ the set of morphisms satisfying the above condition in $D^{b}(A)$. Then the homomorphism space $\mathrm{Hom}_{D_{sg}(A)}(M^{*},N^{*})$ of $D_{sg}(A)$ is the localization of $\mathrm{Hom}_{D^{b}(A)}(M^{*},N^{*})$ over $\mathcal{S}$.
Definition 2.2.
Let $A$ and $B$ be two associative algebras. $A$ and $B$ are called singular equivalent if there is a triangle equivalence functor
$D_{sg}(A)\xrightarrow{\sim}D_{sg}(B).$
There are many examples of singular equivalence. For example, Chen and Sun introduced the singular equivalence of Morita type in [8]. More generally, Wang introduced singular equivalence of Morita type with level in [23]. We give some examples of singular equivalence.
Example 2.3.
Let $S:=k[x_{1},\cdots,x_{m}]$ for some integer $m$ and $1\neq f\in S$ be a nontrivial element. Consider two algebras $S[u]/(f)$ and $S[u,v]/(f+uv)$ by adding variables $u$ and $u,v$ respectively. There are two algebra morphisms
$S\hookrightarrow S[u]/(f)$
given by the natural injection, and
$S[u,v]/(f+uv)\twoheadrightarrow S[u]/(f)$
mapping $v$ to zero. Notice that $S[u]/(f)$ is an $\big{(}S[u,v]/(f+uv)\big{)}\otimes S^{op}$-module.
When we view $S[u]/(f)$ as an $S[u,v]/(f+uv)$-module, it admits a projective resolution as follows:
$0\rightarrow S[u,v]/(f+uv)\xrightarrow{v\cdot}S[u,v]/(f+uv)\rightarrow S[u]/(f)\rightarrow 0$
It follows that $S[u]/(f)\in\mathrm{Perf}\big{(}S[u,v]/(f+uv)\big{)}$. Thus, there is a triangulated functor
$D_{sg}\big{(}S/(f)\big{)}\xrightarrow{(-)\otimes^{\mathbb{L}}_{S/(f)}S[u]/(f)}D_{sg}\big{(}S[u,v]/(f+uv)\big{)}.$
In [16], Knorrer proved that the above triangulated functor is a singular equivalence between $S/(f)$ and $S[u,v]/(f+uv)$.
Example 2.4.
Let $R$ and $S$ be two commutative Noetherian rings, where $R$ is homologically smooth. Let $N$ be an $S\otimes R^{op}$-module such that it is projective over $S$. Set
$A:=\left(\begin{matrix}R&N\\ 0&S\end{matrix}\right),$
and $M:=\left(\begin{matrix}N\\ S\end{matrix}\right)$. Then $M$ is an $S\otimes A^{op}$-module and a projective $S$-module. In [7, Theorem 4.1], Chen showed that the triangulated functor
$D_{sg}(A)\xrightarrow{(-)\otimes^{\mathbb{L}}_{A}M}D_{sg}(S).$
is a singular equivalence. Furthermore, $M$ is also a projective $A^{op}$-module. The triangulated functor
$D_{sg}(S)\xrightarrow{(-)\otimes^{\mathbb{L}}_{S}M^{T}}D_{sg}(A)$
is a singular equivalence, where $M^{T}$ is the transpose of $M$.
2.2. DG category and its DG enhancement
Definition 2.5.
A DG category is a $k$-linear category $\mathcal{D}$ such that the Hom-set $\mathrm{Hom}_{\mathcal{D}}(X,Y)$ consists of complexes of vector spaces, and the composition
$\mathrm{Hom}_{\mathcal{D}}(Y,Z)\otimes\mathrm{Hom}_{\mathcal{D}}(X,Y)\rightarrow\mathrm{Hom}_{\mathcal{D}}(X,Z),$
and the identity morphisms are closed in degree zero.
A DG functor $F:\mathcal{D}\rightarrow\mathcal{D}^{\prime}$ between DG categories is a functor such that the map
$F:\mathrm{Hom}_{\mathcal{D}}(X,Y)\rightarrow\mathrm{Hom}_{\mathcal{D}^{\prime}}\big{(}F(X),F(Y)\big{)}$
is a morphism of complexes.
A DG algebra is a DG category with one object. A homomorphism between DG algebras can be viewed as a DG functor between the corresponding DG categories.
For any DG category $\mathcal{D}$, we can construct a $k$-linear graded category $\mathrm{H}(\mathcal{D})$ given as follows:
  1. (i)
    $\mathrm{Obj}\big{(}\mathrm{H}(\mathcal{D})\big{)}=\mathrm{Obj}(\mathcal{D})$;
  2. (ii)
    $\mathrm{Hom}_{\mathrm{H}(\mathcal{D})}(X,Y)=\mathrm{H}^{\ast}\big{(}\mathrm{Hom}_{\mathcal{D}}(X,Y)\big{)}=:\bigoplus\limits_{i\in\mathbb{Z}}\mathrm{H}^{i}\big{(}\mathrm{Hom}_{\mathcal{D}}(X,Y)\big{)}$, for any objects $X,Y$.
Furthermore, let category $\mathrm{H}^{0}(\mathcal{D})$ be such that
  1. (i)
    $\mathrm{Obj}\big{(}\mathrm{H}^{0}(\mathcal{D})\big{)}=\mathrm{Obj}\big{(}\mathcal{D}\big{)}$;
  2. (ii)
    $\mathrm{Hom}_{\mathrm{H}^{0}(\mathcal{D})}(X,Y)=\mathrm{H}^{0}\big{(}\mathrm{Hom}_{\mathcal{D}}(X,Y)\big{)}$.
Naturally, any DG functor $F$ induces two functors $\mathrm{H}(F)$ and $\mathrm{H}^{0}(F)$.
Definition 2.6.
A DG functor $F:\mathcal{D}\rightarrow\mathcal{D}^{\prime}$ is called a quasi-equivalence if
$F:\mathrm{Hom}_{\mathcal{D}}(X,Y)\rightarrow\mathrm{Hom}_{\mathcal{D}^{\prime}}\big{(}F(X),F(Y)\big{)}$
is a quasi-isomorphism, for any $X,Y\in\mathrm{Obj}(\mathcal{D})$, and $\mathrm{H}^{0}(F)$ is an equivalence.
Now, let $\mathrm{DGCat}$ be the category whose objects are small DG categories and whose morphisms are DG functors. Consider the localization, denoted by $\mathrm{Hqe}$, of $\mathrm{DGCat}$ with respect to quasi-equivalences. We call any morphism in $\mathrm{Hqe}$ a quasi-functor.
For more properties of DG categories such as DG modules of DG categories and tensor functors between them, we refer to the papers [14] and [17].
Given a small DG category $\mathcal{C}$, let $\mathrm{DGMod}(\mathcal{C})$ be the DG category of $\mathcal{C}$-modules, which is defined to be the set $\mathrm{Hom}(\mathcal{C}^{op},\mathcal{D}(k))$ of DG functors, where $\mathcal{D}(k)$ is the canonical DG category of complexes of $k$-linear vector spaces. A $\mathcal{C}$-module is representable if it is contained in the essential image of the Yoneda DG functor
$\mathrm{Y}_{dg}^{\mathcal{C}}:\mathcal{C}\rightarrow\mathrm{DGMod}(\mathcal{C}),\,\,\,X\mapsto\mathrm{Hom}_{\mathcal{C}}(-,X).$
Definition 2.7.
A DG category $\mathcal{C}$ is called pretriangulated if the essential image of the functor
$\mathrm{H}^{0}\big{(}\mathrm{Y}_{dg}^{\mathcal{C}}\big{)}:\mathrm{H}^{0}(\mathcal{C})\rightarrow\mathrm{H}^{0}\big{(}\mathrm{DGMod}(\mathcal{C})\big{)}$
is a triangulated subcategory.
Definition 2.8.
Let $\mathrm{D}$ be a triangulated category.
(1) A DG enhancement of $\mathrm{D}$ is a pair $(\mathcal{D},\psi)$, where $\mathcal{D}$ is a pre-triangulated DG category and $\psi:\mathrm{H}^{0}(\mathcal{D})\rightarrow\mathrm{D}$ is an exact equivalence.
(2) $\mathrm{D}$ has a unique DG enhancement if for two DG enhancements $(\mathcal{D},\psi)$ and $(\mathcal{D}^{\prime},\psi^{\prime})$, there exists a quasi-functor $F:\mathcal{D}\rightarrow\mathcal{D}^{\prime}$ such that $\mathrm{H}^{0}(F)$ is an equivalence.
In the definition, the equivalence of $\mathrm{H}^{0}(F)$ means that $F$ is an equivalence in $\mathrm{Hqe}$.
3. The Tate-Hochschild cohomology
3.1. Generalized Tate-Hochschild complex of associative algebra
The notion of Tate-Hochschild cohomology was introduced by Buchweitz (see [5]).
Definition 3.1.
Let $\Lambda$ be an associative algebra. Its $i$-th Tate-Hochschild cohomology, denoted by $\mathrm{HH}_{sg}^{i}(\Lambda)$, is $\mathrm{Hom}_{D_{sg}(\Lambda^{e})}(\Lambda,\Lambda[i]),$ where $\Lambda^{e}:=\Lambda\otimes\Lambda^{op}$.
Later, for any associative algebra $\Lambda$, Wang defined the singular Hochschild complex whose cohomologies are the Tate-Hochschild cohomologies ([24]). This complex is constructed as the colimit of a sequence of complexes. He also constructed a complex, called the generalized Tate-Hochschild complex of $\Lambda$:
[…]
whose cohomologies, in the case of $\Lambda$ being self-injective, are exactly the Tate-Hochschild cohomologies, where $\Lambda^{\vee}:=\mathrm{Hom}_{\Lambda^{e}}(\Lambda,\Lambda\otimes\Lambda)$, the differential $\mu$ is given by the multiplication of $\Lambda$: $\mu(x\otimes y)=xy$, $C_{\ast}(\Lambda,\Lambda^{\vee})$ is the Hochschild chain complex of $\Lambda$ for $\Lambda^{e}$-module $\Lambda^{\vee}$, and $C^{\ast}(\Lambda,\Lambda)$ is the Hochschild cochain complex of $\Lambda$.
In the following, we introduce a new complex for an arbitrary associative algebra $A$ such that its cohomologies are also the Tate-Hochschild cohomologies (see Proposition 3.6 below). The underlying vector space of this complex is described explicitly. Moreover, when $A$ is self-injective, this complex coincides with $\mathcal{D}^{\ast}(A,A)$.
Let us first recall the DG quotient of a DG category, introduced by Drinfeld (see [9]).
Definition 3.2.
Let $\mathcal{A}$ be a DG category and $\mathcal{B}\subset\mathcal{A}$ be a full DG subcategory of $\mathcal{A}$. A DG quotient of $\mathcal{A}$ by $\mathcal{B}$ is a diagram consisting of DG categories and DG functors
[…]
satisfying that
  1. (1)
    the above DG functor $\tilde{\mathcal{A}}\xrightarrow{\sim}\mathcal{A}$ is a quasi-equivalence;
  2. (2)
    the functor $\mathrm{H}(\pi):\mathrm{H}(\tilde{\mathcal{A}})\rightarrow\mathrm{H}(\mathcal{C})$ between corresponding homotopy categories induced by $\pi$ is essentially surjective;
  3. (3)
    $\mathrm{H}(\pi)$ gives a triangulated functor $\tilde{\mathcal{A}}^{tr}\rightarrow\mathcal{C}^{tr}$, which induces an equivalence:
    $\mathcal{A}^{tr}/\mathcal{B}^{tr}\xrightarrow{\sim}\mathcal{C}^{tr},$
    where $\mathcal{A}^{tr}/\mathcal{B}^{tr}$ is the Verdier quotient of $\mathcal{A}^{tr}$ by $\mathcal{B}^{tr}$, $\mathcal{A}^{tr}$ (resp. $\tilde{\mathcal{A}}^{tr},\mathcal{B}^{tr},\mathcal{C}^{tr}$) represents the triangulated category $\mathrm{H}(\mathcal{A}^{pretr})$ (resp. $\mathrm{H}(\tilde{\mathcal{A}}^{pretr}),\mathrm{H}(\mathcal{B}^{pretr}),\mathrm{H}(\mathcal{C}^{pretr})$), and $\mathcal{A}^{pretr}$ (resp. $\tilde{\mathcal{A}}^{pretr},\mathcal{B}^{pretr},\mathcal{C}^{pretr}$) is a certain pre-triangulated DG category on $\mathcal{A}$ (resp. $\tilde{\mathcal{A}},\mathcal{B},\mathcal{C}$).
From [9, §3], we know that, for any DG category $\mathcal{A}$ and its full DG subcategory $\mathcal{B}\subset\mathcal{A}$, the DG quotient of $\mathcal{A}$ by $\mathcal{B}$ exists. On the other hand, it is well known that both $D^{b}(A)$ and $\mathrm{Perf}(A)$ admit a canonical DG category structure. Here, we denote by $\mathcal{P}\mathrm{erf}(A)$ and $\mathcal{D}^{b}(A)$ the DG categories corresponding to $\mathrm{Perf}(A)$ and $D^{b}(A)$ respectively. Moreover, both $\mathcal{P}\mathrm{erf}(A)$ and $\mathcal{D}^{b}(A)$ have canonical pre-triangulated structures. Hence, by taking the DG quotient, $D_{sg}(A)$ is endowed with a DG category structure which is a DG enhancement for $D_{sg}(A)$. We denote by $S_{dg}(A)$ this DG category associated to $D_{sg}(A)$.
In [15, Theorem 1.1], Keller realized the Tate-Hochschild cohomologies of an algebra $A$ as the Hochschild cohomologies of $S_{dg}(A)$ in the following theorem.
Theorem 3.3.
There is a canonical isomorphism of graded algebras between the Tate-Hochschild cohomologies $\mathrm{HH}^{\ast}_{sg}(A)$ of $A$ and the Hochschild cohomologies of $S_{dg}(A)$.
Next we recall the following proposition (see [9, 1.3.1]).
Proposition 3.4.
Let $M^{\ast},N^{\ast}\in\mathrm{Obj}\big{(}S_{dg}(A)\big{)}$. Then
$\mathrm{Hom}_{S_{dg}(A)}(M^{\ast},N^{\ast})\cong\mathrm{Cone}\big{(}h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)}^{\mathbb{L}}\tilde{h}_{M^{\ast}}\rightarrow\mathrm{Hom}_{\mathcal{D}^{b}(A)}(M^{\ast},N^{\ast})\big{)}\quad\textup{in}\,\,D(k),$
where $h_{N^{\ast}}(-):=\mathrm{Hom}_{\mathcal{D}^{b}(A)}(-,N^{\ast})$ as a $\mathcal{P}\mathrm{erf}(A)$-module, $\tilde{h}_{M^{\ast}}(-):=\mathrm{Hom}_{\mathcal{D}^{b}(A)}(M^{\ast},-)$ as a $\mathcal{P}\mathrm{erf}(A)^{op}$-module and $\mathcal{P}\mathrm{erf}(A)^{op}$ is the opposite DG category of $\mathcal{P}\mathrm{erf}(A)$.
In above proposition, the morphism $h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)}^{\mathbb{L}}\tilde{h}_{M^{\ast}}\rightarrow\mathrm{Hom}_{\mathcal{D}^{b}(A)}(M^{\ast},N^{\ast})$ in above proposition is given by composition of morphisms in $\mathcal{D}^{b}(A)$. In the meantime, we have the following.
Lemma 3.5.
With the setting of the above proposition,
$h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)}^{\mathbb{L}}\tilde{h}_{M^{\ast}}\cong N^{\ast}\otimes^{\mathbb{L}}_{A}(M^{\ast})^{\vee^{\mathbb{L}}}$
in $D(k)$, where $(M^{\ast})^{\vee^{\mathbb{L}}}:=\mathbb{R}\mathrm{Hom}_{A}(M^{\ast},A)$ is an object of $D^{b}(A^{op})$.
Proof.
Consider the natural inclusion functor of DG categories
$\Psi_{A}:A^{op}\rightarrow\mathcal{P}\mathrm{erf}(A),$
sending the unique object of $A^{op}$ to $A$-module complex $A$. By the result of Keller in [14, Theorem 8.1], the triangulated functor
$(\Psi_{A})_{*}:D\big{(}\mathcal{P}\mathrm{erf}(A)\big{)}\rightarrow D(A^{op})$
induced by $\Psi_{A}$, is a triangle equivalence. In the same way, we get the equivalence
$(\Psi^{op}_{A})_{*}:D\big{(}\mathcal{P}\mathrm{erf}(A)^{op}\big{)}\rightarrow D(A).$
By the definition of tensor product of DG modules over a DG category (see [17, 3.5]), we know that
$h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)}\tilde{h}_{M^{\ast}}$
$\cong\mathrm{Cone}\Big{(}\bigoplus_{X,Y\in\mathcal{P}\mathrm{erf}(A)}h_{N^{\ast}}(X)\otimes_{k}\mathrm{Hom}_{\mathcal{P}\mathrm{erf}(A)}(Y,X)\otimes_{k}\tilde{h}_{M^{\ast}}(Y)$
$\quad\quad\quad\quad\quad\quad\rightarrow\bigoplus_{X\in\mathcal{P}\mathrm{erf}(A)}h_{N^{\ast}}(X)\otimes_{k}\tilde{h}_{M^{\ast}}(X)\Big{)}$
$\cong\mathrm{Cone}\Big{(}\bigoplus_{X,Y\in\mathcal{P}\mathrm{erf}(A)}\mathrm{Hom}_{\mathcal{D}^{b}(A)}(X,N^{\ast})\otimes_{k}\mathrm{Hom}_{\mathcal{P}\mathrm{erf}(A)}(Y,X)\otimes_{k}\mathrm{Hom}_{\mathcal{D}^{b}(A)}(M^{\ast},Y)$
$\quad\quad\quad\quad\quad\quad\rightarrow\bigoplus_{X\in\mathcal{P}\mathrm{erf}(A)}\mathrm{Hom}_{\mathcal{D}^{b}(A)}(X,N^{\ast})\otimes_{k}\mathrm{Hom}_{\mathcal{D}^{b}(A)}(M^{\ast},X)\Big{)}.$
If we replace $N^{\ast}$ by its projective $A$-module resolution, say, $P_{N^{*}}$, then by [17, 3.5], $h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)}^{\mathbb{L}}\tilde{h}_{M^{\ast}}$ is isomorphic to $h_{P_{N^{\ast}}}\otimes_{\mathcal{P}\mathrm{erf}(A)}\tilde{h}_{M^{\ast}}$. It follows that
$h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)}^{\mathbb{L}}\tilde{h}_{M^{\ast}}$
$\cong\mathrm{Cone}\Big{(}\bigoplus_{X,Y\in\mathcal{P}\mathrm{erf}(A)}h_{P_{N^{\ast}}}(X)\otimes_{k}\mathrm{Hom}_{\mathcal{P}\mathrm{erf}(A)}(Y,X)\otimes_{k}\tilde{h}_{M^{\ast}}(Y)$
$\quad\quad\quad\quad\quad\quad\rightarrow\bigoplus_{X\in\mathcal{P}\mathrm{erf}(A)}h_{P_{N^{\ast}}}(X)\otimes_{k}\tilde{h}_{M^{\ast}}(X)\Big{)}$
$\cong\mathrm{Cone}\Big{(}\bigoplus_{X,Y\in\mathcal{P}\mathrm{erf}(A)}\mathrm{Hom}_{\mathcal{D}^{b}(A)}(X,P_{N^{\ast}})\otimes_{k}\mathrm{Hom}_{\mathcal{P}\mathrm{erf}(A)}(Y,X)\otimes_{k}\mathrm{Hom}_{\mathcal{D}^{b}(A)}(M^{\ast},Y)$
$\quad\quad\quad\quad\quad\quad\rightarrow\bigoplus_{X\in\mathcal{P}\mathrm{erf}(A)}\mathrm{Hom}_{\mathcal{D}^{b}(A)}(X,P_{N^{\ast}})\otimes_{k}\mathrm{Hom}_{\mathcal{D}^{b}(A)}(M^{\ast},X)\Big{)}.$
On the other hand, we know
$N^{\ast}\otimes^{\mathbb{L}}_{A}(M^{\ast})^{\vee^{\mathbb{L}}}$ $\cong P_{N^{\ast}}\otimes_{A}(M^{\ast})^{\vee^{\mathbb{L}}}$
$\cong\mathrm{Hom}_{\mathcal{D}^{b}(A)}(A,P_{N^{\ast}})\otimes_{A}\mathrm{Hom}_{\mathcal{D}^{b}(A)}(M^{\ast},A)$
$\cong\mathrm{Cone}\Big{(}\mathrm{Hom}_{\mathcal{D}^{b}(A)}(A,P_{N^{\ast}})\otimes\mathrm{Hom}_{\mathcal{D}^{b}(A)}(A,A)\otimes\mathrm{Hom}_{\mathcal{D}^{b}(A)}(M^{\ast},A)$
$\quad\quad\quad\quad\quad\quad\rightarrow\mathrm{Hom}_{\mathcal{D}^{b}(A)}(A,P_{N^{\ast}})\otimes\mathrm{Hom}_{\mathcal{D}^{b}(A)}(M^{\ast},A)\Big{)}.$
Hence, we obtain a morphism between complexes
$\eta_{N^{*}}^{M^{*}}:N^{\ast}\otimes^{\mathbb{L}}_{A}(M^{\ast})^{\vee^{\mathbb{L}}}\hookrightarrow h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)}^{\mathbb{L}}\tilde{h}_{M^{\ast}}$
given by
$\mathrm{Cone}\Big{(}\mathrm{Hom}_{D^{b}(A)}(A,P_{N^{\ast}})\otimes\mathrm{Hom}_{D^{b}(A)}(A,A)\otimes\mathrm{Hom}_{D^{b}(A)}(M^{\ast},A)$
$\quad\quad\quad\quad\quad\quad\rightarrow\mathrm{Hom}_{D^{b}(A)}(A,P_{N^{\ast}})\otimes\mathrm{Hom}_{D^{b}(A)}(M^{\ast},A)\Big{)}$
$\hookrightarrow$ $\mathrm{Cone}\Big{(}\bigoplus_{X,Y\in\mathcal{P}\mathrm{erf}(A)}\mathrm{Hom}_{\mathcal{D}^{b}(A)}(X,P_{N^{\ast}})\otimes_{k}\mathrm{Hom}_{\mathcal{P}\mathrm{erf}(A)}(Y,X)\otimes_{k}\mathrm{Hom}_{\mathcal{D}^{b}(A)}(M^{\ast},Y)$
$\quad\quad\quad\quad\quad\quad\rightarrow\bigoplus_{X\in\mathcal{P}\mathrm{erf}(A)}\mathrm{Hom}_{\mathcal{D}^{b}(A)}(X,P_{N^{\ast}})\otimes_{k}\mathrm{Hom}_{\mathcal{D}^{b}(A)}(M^{\ast},X)\Big{)}.$
Obviously, $\eta_{N^{*}}^{M^{*}}$ is a quasi-isomorphism if and only if the induced morphism
$\mathbb{R}\mathrm{Hom}_{k}(\eta_{N^{*}}^{M^{*}},k):\mathbb{R}\mathrm{Hom}_{k}\big{(}h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)}^{\mathbb{L}}\tilde{h}_{M^{\ast}},k\big{)}\twoheadrightarrow\mathbb{R}\mathrm{Hom}_{k}\big{(}N^{\ast}\otimes^{\mathbb{L}}_{A}(M^{\ast})^{\vee^{\mathbb{L}}},k\big{)}$
is a quasi-isomorphism.
Now, by adjunction of functors, we obtain that
$\mathbb{R}\mathrm{Hom}_{k}\big{(}h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)^{op}}^{\mathbb{L}}\tilde{h}_{M^{\ast}},k\big{)}\cong\mathrm{Hom}_{\mathcal{P}\mathrm{erf}(A)}\big{(}h_{N^{\ast}},\mathbb{D}(\tilde{h}_{M^{\ast}})\big{)}$
and
$\mathbb{R}\mathrm{Hom}_{k}\big{(}N^{\ast}\otimes^{\mathbb{L}}_{A}(M^{\ast})^{\vee^{\mathbb{L}}},k\big{)}\cong\mathbb{R}\mathrm{Hom}_{A}\big{(}N^{*},\mathbb{D}((M^{\ast})^{\vee^{\mathbb{L}}})\big{)},$
where $\mathbb{D}(-):=\mathbb{R}\mathrm{Hom}_{k}(-,k)$. Note that, for any object $X$ of $\mathcal{P}\mathrm{erf}(A)$, $\mathbb{D}(\tilde{h}_{M^{\ast}})(X):=\mathbb{D}\big{(}\tilde{h}_{M^{\ast}}(X)\big{)}$. Meanwhile, we have
$(\Psi^{op}_{A})_{*}(h_{N^{*}})\cong N^{*}\quad\textup{ and}\quad(\Psi^{op}_{A})_{*}(\tilde{h}_{M^{*}})\cong(M^{\ast})^{\vee^{\mathbb{L}}}.$
Hence,
$(\Psi^{op}_{A})_{*}\big{(}\mathbb{D}(\tilde{h}_{M^{*}})\big{)}\cong\mathbb{D}\big{(}(M^{\ast})^{\vee^{\mathbb{L}}}\big{)}.$
Since $\Psi^{op}_{A}$ is a triangle equivalence, we have that $\mathrm{Hom}_{\mathcal{P}\mathrm{erf}(A)^{op}}\big{(}h_{N^{\ast}},\mathbb{D}(\tilde{h}_{M^{\ast}})\big{)}$ is isomorphic to $\mathbb{R}\mathrm{Hom}_{A}\big{(}N^{*},\mathbb{D}((M^{\ast})^{\vee^{\mathbb{L}}})\big{)}$ in $D(k)$. Thus, $\mathbb{R}\mathrm{Hom}_{k}(\eta_{N^{*}}^{M^{*}},k)$ is a quasi-isomorphism. It implies that
$h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)}^{\mathbb{L}}\tilde{h}_{M^{\ast}}\cong N^{\ast}\otimes^{\mathbb{L}}_{A}(M^{\ast})^{\vee^{\mathbb{L}}}$
in $D(k)$. We thus completed the proof. ∎
Combining Proposition 3.4 with Lemma 3.5, we obtain that
$\mathrm{Hom}_{S_{dg}(A)}(M^{\ast},N^{\ast})\cong\mathrm{Cone}\big{(}N^{\ast}\otimes^{\mathbb{L}}_{A}(M^{\ast})^{\vee^{\mathbb{L}}}\rightarrow\mathrm{Hom}_{\mathcal{D}^{b}(A)}(M^{\ast},N^{\ast})\big{)}$
in $D(k)$. Now let $R$ be a commutative Noetherian ring, and in the above isomorphism set $A=R^{e}$ and $M^{\ast}=N^{\ast}=R$ as $R\otimes R^{op}$-module. Then we get
$\mathrm{HH}^{i}_{sg}(R):=\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])=\mathrm{H}^{i}\big{(}\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)\big{)}$
$\cong\mathrm{H}^{i}\left(\mathrm{Cone}\big{(}R\otimes^{\mathbb{L}}_{R^{e}}(R)^{\vee^{\mathbb{L}}}\rightarrow\mathrm{Hom}_{\mathcal{D}^{b}(R^{e})}(R,R)\big{)}\right)$
for any $i\in\mathbb{Z}$.
Now, we consider the Bar resolution $\mathrm{Bar}(R)\twoheadrightarrow R$ over $R^{e}$. There is a double complex:
[…]
given by $\mathrm{Bar}(R)\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$. View this double complex as a cochain complex, and denote it by $E^{\ast}$. Also, denote by $F^{\ast}$ the cochain complex obtained from the double complex $\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),\mathrm{Bar}(R))$. We get a complex $\mathrm{Cone}(E^{\ast}\rightarrow F^{\ast})$ and obtain the following proposition for this complex.
Proposition 3.6.
With the above settings,
$\mathrm{HH}^{i}_{sg}(R)\cong\mathrm{H}^{i}\big{(}\mathrm{Cone}(E^{\ast}\rightarrow F^{\ast})\big{)}$
holds for any $i\in\mathbb{Z}$.
3.2. Hochschild cohomology of singularity category
In this subsection, we give the proofs of the statements in the steps (1) and (2) in §1.3.
Our goal now is to endow $\mathrm{Cone}(E^{\ast}\rightarrow F^{*})$ with a natural multiplication. First, by the construction of $F^{\ast}=\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),\mathrm{Bar}(R))$, $F^{\ast}$ has a natural multiplication structure from the endomorphism ring. Second, the following map
$E^{\ast}\otimes E^{\ast}\cong$ $\Big{(}\mathrm{Bar}(R)\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}\Big{)}\otimes\Big{(}\mathrm{Bar}(R)\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}\Big{)}$
$\cong\mathrm{Bar}(R)\otimes_{R^{e}}\Big{(}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}\otimes\mathrm{Bar}(R)\Big{)}\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
$\xrightarrow{invol.}\mathrm{Bar}(R)\otimes_{R^{e}}R^{e}\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
$\cong\mathrm{Bar}(R)\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}=E^{\ast},$
induced by involution, gives a natural multiplication structure on $E^{\ast}$. Finally, there is a canonical $\Big{(}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),\mathrm{Bar}(R)\big{)}\Big{)}^{e}$-module structure on $\mathrm{Bar}(R)\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$. Therefore, viewing $\mathrm{Cone}(E^{\ast}\rightarrow F^{*})$ as a semi-product $F^{\ast}\ltimes E^{\ast}$, we obtain on $\mathrm{Cone}(E^{\ast}\rightarrow F^{*})$ a natural multiplication structure.
Next, recall that the canonical morphism $E^{\ast}\rightarrow F^{\ast}$ of complexes is given by the following composition map:
$\mathrm{Bar}(R)\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$ $\cong\mathrm{Hom}_{R^{e}}\big{(}R^{e},\mathrm{Bar}(R)\big{)}\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
$\rightarrow\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),\mathrm{Bar}(R)).$
This morphism together with differentials on $F^{\ast}$ and $E^{\ast}$ gives the natural differential on $\mathrm{Cone}(E^{\ast}\rightarrow F^{*})$. It is easy to check that the differential is compatible with the multiplication on $\mathrm{Cone}(E^{\ast}\rightarrow F^{*})$. Thus, $\mathrm{Cone}(E^{\ast}\rightarrow F^{*})$ is a differential graded algebra.
We next recall the support schemes of object in $S_{sg}(R)$ and of $S_{sg}(R)$.
Definition 3.7.
(1) Let $Q^{\ast}$ be an object in $S_{sg}(R)$. The support scheme $\mathrm{Supp}(Q^{\ast})$ of $Q^{\ast}$ is a subscheme of $\mathrm{Spec}(R)$:
$\mathrm{Supp}(Q^{\ast}):=\big{\{}\mathbf{p}\in\mathrm{Spec}(R)\mid Q^{\ast}_{\mathbf{p}}\ncong 0\,\textup{in}\,D_{sg}(R_{\mathbf{p}})\big{\}}.$
(2) The support scheme $\mathrm{Supp}(S_{dg}(R))$ of $S_{dg}(R)$ is a subscheme of $\mathrm{Spec}(R)$ :
$\mathrm{Supp}(S_{dg}(R)):=\bigcap\limits_{Q^{\ast}\in\mathrm{Obj}\big{(}S_{dg}(R)\big{)}}\mathrm{Supp}(Q^{\ast}).$
We have the following.
Proposition 3.8.
Let $R$ be a commutative Noetherian ring. Let $R/I$ be the coordinate ring of singular locus on $\mathrm{Spec}(R)$. Then there is an algebra injection
$\iota_{I}:R/I\hookrightarrow\mathrm{HH}^{0}_{sg}(R).$
Proof.
First, there is a canonical DG algebra morphism
$F^{\ast}\hookrightarrow\mathrm{Cone}(E^{\ast}\rightarrow F^{*}),$
which induces the following algebra homomorphism
$\pi_{R}:R\cong\mathrm{HH}^{0}(R)\cong\mathrm{H}^{0}(F^{*})\rightarrow\mathrm{H}^{0}\big{(}\mathrm{Cone}(E^{\ast}\rightarrow F^{*})\big{)}\cong\mathrm{HH}_{sg}^{0}(R).$
Next, the composition of algebra morphisms:
$R\xrightarrow{\pi_{R}}\mathrm{HH}_{sg}^{0}(R)\hookrightarrow\bigoplus_{i}\mathrm{HH}_{sg}^{i}(R)\cong\bigoplus_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$
gives the canonical $R$-module structure on $\bigoplus\limits_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$.
Since $\bigoplus\limits_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$ contains the unit $\mathrm{id}$ and $\pi_{R}(1_{R})=\mathrm{id}$, its annihilator, as an $R$-module, is trivial. Hence, when we view $\bigoplus\limits_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$ as a sheaf on $\mathrm{Spec}(R)$, to prove the proposition, it suffices to show that the support scheme of $\bigoplus\limits_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$ is $V(I)\subseteq\mathrm{Spec}(R)$, which is equivalent to showing that the support scheme of $\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)$ in $D(k)$ is exactly $V(I)$.
Let $\mathbf{p}$ be a point in $\mathrm{Spec}(R)$. We know that
$\left(\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)\right)_{\mathbf{p}}$ $\cong\mathrm{Cone}(E^{\ast}\rightarrow F^{*})_{\mathbf{p}}$
$\cong\mathrm{Cone}(E^{\ast}_{\mathbf{p}}\rightarrow F^{*}_{\mathbf{p}})$
in $D(k)$. On the one hand,
$F^{*}_{\mathbf{p}}$ $=\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),R)_{\mathbf{p}}$
$\cong\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),R_{\mathbf{p}})$
$\cong\mathbb{R}\mathrm{Hom}_{R_{\mathbf{p}}^{e}}(\mathrm{Bar}(R)\otimes_{R^{e}}R_{\mathbf{p}}^{e},R_{\mathbf{p}})$
$\cong\mathbb{R}\mathrm{Hom}_{R_{\mathbf{p}}^{e}}(\mathrm{Bar}(R_{\mathbf{p}}),R_{\mathbf{p}})$
holds in $D(k)$, where $\mathrm{Bar}(R)\otimes_{R^{e}}R_{\mathbf{p}}^{e}$ is also a projective $R_{\mathbf{p}}^{e}$-module resolution of $R_{\mathbf{p}}$. On the other hand,
$E^{*}_{\mathbf{p}}$ $=\left(\mathrm{Bar}(R)\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}\right)_{\mathbf{p}}$
$\cong\left(R\otimes_{R^{e}}\overline{\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}}\right)_{\mathbf{p}}$
$\cong R_{\mathbf{p}}\otimes_{R^{e}}\overline{\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}}$
$\cong R_{\mathbf{p}}\otimes_{R_{\mathbf{p}}^{e}}R_{\mathbf{p}}^{e}\otimes_{R^{e}}\overline{\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}}$
$\cong(R_{\mathbf{p}}\otimes_{R_{\mathbf{p}}^{e}}R_{\mathbf{p}}^{e})\otimes^{\mathbb{L}}_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
$\cong(R_{\mathbf{p}}\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}R_{\mathbf{p}}^{e})\otimes^{\mathbb{L}}_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
$\cong\mathrm{Bar}(R_{\mathbf{p}})\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}R_{\mathbf{p}}^{e}\otimes^{\mathbb{L}}_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
$\cong\mathrm{Bar}(R_{\mathbf{p}})\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}\Big{(}R_{\mathbf{p}}^{e}\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}\Big{)}$
$\cong\mathrm{Bar}(R_{\mathbf{p}})\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R_{\mathbf{p}}^{e}\big{)}$
$\cong\mathrm{Bar}(R_{\mathbf{p}})\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}\mathrm{Hom}_{R_{\mathbf{p}}^{e}}\big{(}\mathrm{Bar}(R)\otimes_{R^{e}}R_{\mathbf{p}}^{e},R_{\mathbf{p}}^{e}\big{)}$
$\cong\mathrm{Bar}(R_{\mathbf{p}})\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}\mathrm{Hom}_{R_{\mathbf{p}}^{e}}\big{(}\mathrm{Bar}(R_{\mathbf{p}}),R_{\mathbf{p}}^{e}\big{)}$
in $D(k)$, where $\overline{\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}}$ is a flat resolution of $\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$ over $R^{e}$. In the meantime, when we replace $R$ by $R_{\mathbf{q}}$, we obtain that
$\mathrm{Cone}(E^{\ast}_{\mathbf{p}}\rightarrow F^{*}_{\mathbf{p}})\cong\mathrm{Hom}_{S_{dg}(R_{\mathbf{p}}^{e})}(R_{\mathbf{p}},R_{\mathbf{p}})$
in $D(k)$ as the argument in §2.1. Thus, we get
$\left(\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)\right)_{\mathbf{p}}\cong\mathrm{Hom}_{S_{dg}(R_{\mathbf{p}}^{e})}(R_{\mathbf{p}},R_{\mathbf{p}})$
as algebras.
Now we assume that $\mathbf{p}$ is a non-singular point. Since ${\mathbf{p}}$ is non-singular, $R_{\mathbf{p}}$ is a homologically smooth algebra. Then $\mathrm{Hom}_{S_{dg}(R_{\mathbf{p}}^{e})}(R_{\mathbf{p}},R_{\mathbf{p}})$ is a trivial in $D(k)$. Hence $\left(\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)\right)_{\mathbf{p}}$ is also a trivial in $D(k)$. Thus, ${\mathbf{p}}$ is not contained in the support scheme $\mathrm{Supp}(S_{dg}(R))$ of $\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)$ in $D(k)$, which implies that
$\mathrm{Supp}(S_{dg}(R))\subseteq V(I):=\mathrm{Spec}(R/I).$
Next, let $\mathbf{q}$ be a singular point in $\mathrm{Spec}(R)$, i.e., $\mathbf{q}\in V(I)$. As in above argument, we know that $\left(\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)\right)_{\mathbf{q}}\cong\mathrm{Hom}_{S_{dg}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}},R_{\mathbf{q}})$.
Assume that $\mathrm{Hom}_{S_{dg}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}},R_{\mathbf{q}})$ is a trivial in $D(k)$. It implies that $\mathrm{HH}^{0}_{sg}(R_{\mathbf{q}})$ is a trivial algebra. From the distinguished triangle
$F^{\ast}_{\mathbf{q}}\rightarrow\mathrm{Cone}(E_{\mathbf{q}}^{\ast}\rightarrow F_{\mathbf{q}}^{*})\rightarrow E_{\mathbf{q}}^{\ast}[1],$
we have the long exact sequence
$\cdots\rightarrow\mathrm{H}^{0}(E_{\mathbf{q}}^{\ast})\rightarrow\mathrm{H}^{0}(F_{\mathbf{q}}^{\ast})\xrightarrow{\pi_{R}}\mathrm{H}^{0}\big{(}\mathrm{Cone}(E_{\mathbf{q}}^{\ast}\rightarrow F_{\mathbf{q}}^{*})\big{)}\rightarrow\mathrm{H}^{1}(E_{\mathbf{q}}^{\ast})\rightarrow\cdots,$
which induces a surjection
$\Theta:\mathrm{H}^{0}(E_{\mathbf{q}}^{\ast})\twoheadrightarrow\mathrm{H}^{0}(F_{\mathbf{q}}^{\ast})$
since
$\mathrm{H}^{0}\big{(}\mathrm{Cone}(E_{\mathbf{q}}^{\ast}\rightarrow F_{\mathbf{q}}^{*})\big{)}=\mathrm{HH}^{0}_{sg}(R_{\mathbf{q}})=0.$
Thus, for the identity element $\mathrm{id}_{\mathrm{Bar}(R_{\mathbf{q}})}\in\mathrm{H}^{0}(F_{\mathbf{q}}^{\ast})$, we have
$\mathrm{id}_{\mathrm{Bar}(R_{\mathbf{q}})}=\Theta\big{(}\sum^{m}_{i=1}\alpha_{i}\otimes_{R_{\mathbf{q}}^{e}}\beta_{i}\big{)}$
in $\mathrm{H}^{0}(F_{\mathbf{q}}^{\ast})$ for some $\alpha_{i}\in\left(\mathbb{R}\mathrm{Hom}_{R_{\mathbf{q}}^{e}}(\mathrm{Bar}(R_{\mathbf{q}}),R_{\mathbf{q}}^{e})\right)^{r_{i}}=\mathrm{Hom}_{R_{\mathbf{q}}^{e}}(\mathrm{Bar}(R_{\mathbf{q}})^{-r_{i}},R_{\mathbf{q}}^{e})$ and $\beta_{i}\in\mathrm{Bar}(R_{\mathbf{q}})^{-r_{i}}$. Here, we use $(-)^{j}$ to denote the degree $j$ component of complex for any $j\in\mathbb{Z}$. For reader’s convenience, we view $\Theta\big{(}\sum^{m}_{i=1}\alpha_{i}\otimes_{R_{\mathbf{q}}^{e}}\beta_{i}\big{)}$ as an element in $F^{\ast}$. It follows that
$\mathrm{id}_{\mathrm{Bar}(R_{\mathbf{q}})}-\Theta\big{(}\sum^{m}_{i=1}\alpha_{i}\otimes_{R_{\mathbf{q}}^{e}}\beta_{i}\big{)}\in\mathrm{Im}(d_{F_{\mathbf{q}}^{\ast}}),$
where $d_{F_{\mathbf{q}}^{\ast}}$ is the differential of $F_{\mathbf{q}}^{\ast}$. Moreover, we have that
$\Theta\big{(}\sum^{m}_{i=1}\alpha_{i}\otimes_{R_{\mathbf{q}}^{e}}\beta_{i}\big{)}\in\bigoplus_{1\leq i\leq m}\mathrm{Hom}_{R_{\mathbf{q}}^{e}}\big{(}\mathrm{Bar}(R_{\mathbf{q}})^{-r_{i}},\mathrm{Bar}(R_{\mathbf{q}})^{-r_{i}}\big{)}.$
Let $l$ be an integer which is greater than all $r_{i}$. Let $N$ be an $R_{\mathbf{q}}^{e}$-module and $f\in\mathrm{Hom}_{D^{b}(R_{\mathbf{q}}^{e})}\big{(}\mathrm{Bar}(R_{\mathbf{q}}),N[l]\big{)}$. Then $f$ is represented by a cycle, denoted by $\tilde{f}\in\mathrm{Hom}_{R_{\mathbf{q}}^{e}}(\mathrm{Bar}(R_{\mathbf{q}})^{-l},N)$, in the complex $\mathbb{R}\mathrm{Hom}_{R_{\mathbf{q}}^{e}}(\mathrm{Bar}(R_{\mathbf{q}}),N)$. Thus we get the following composition
$\tilde{f}\circ\Theta\big{(}\sum^{m}_{i=1}\alpha_{i}\otimes_{R_{\mathbf{q}}^{e}}\beta_{i}\big{)}=0\in\mathbb{R}\mathrm{Hom}_{R_{\mathbf{q}}^{e}}(\mathrm{Bar}(R_{\mathbf{q}}),N)$
since $l\neq r_{i}$ for any $i$. In the meantime, we have that
$\tilde{f}\circ\Big{(}\mathrm{id}_{\mathrm{Bar}(R_{\mathbf{q}})}-\Theta\big{(}\sum^{m}_{i=1}\alpha_{i}\otimes_{R_{\mathbf{q}}^{e}}\beta_{i}\big{)}\Big{)}$
is in the image of the differential of $\mathbb{R}\mathrm{Hom}_{R_{\mathbf{q}}^{e}}(\mathrm{Bar}(R_{\mathbf{q}}),N)$ since
$\mathrm{id}_{\mathrm{Bar}(R_{\mathbf{q}})}-\Theta\big{(}\sum^{m}_{i=1}\alpha_{i}\otimes_{R_{\mathbf{q}}^{e}}\beta_{i}\big{)}\in\mathrm{Im}(d_{F_{\mathbf{q}}^{\ast}}).$
It implies that $\tilde{f}$ is in the image of the differential of $\mathbb{R}\mathrm{Hom}_{R_{\mathbf{q}}^{e}}(\mathrm{Bar}(R_{\mathbf{q}}),N)$. Thus, we obtain that $f=0$ in $\mathrm{Hom}_{D^{b}(R_{\mathbf{q}}^{e})}\big{(}\mathrm{Bar}(R_{\mathbf{q}}),N[l]\big{)}$. It follows that
$\mathrm{Hom}_{D^{b}(R_{\mathbf{q}}^{e})}\big{(}\mathrm{Bar}(R_{\mathbf{q}}),N[l]\big{)}\cong\mathrm{Hom}_{D^{b}(R_{\mathbf{q}}^{e})}\big{(}R_{\mathbf{q}},N[l]\big{)}=0.$
The projective dimension of $R_{\mathbf{q}}$, as an $R^{e}_{\mathbf{q}}$-module, is less that $l$. It suggests that $R_{\mathbf{q}}$ is homologically smooth. Thus we obtain that $\mathbf{q}$ is a non-singular point in $\mathrm{Spec}(R)$. It is a contradiction to the fact that $\mathbf{q}$ is singular in $\mathrm{Spec}(R)$. Hence the assumption that
$\mathrm{Hom}_{S_{dg}(R_{\mathbf{q}}^{e})}\big{(}\mathrm{Bar}(R_{\mathbf{q}}),\mathrm{Bar}(R_{\mathbf{q}})\big{)}\cong\mathrm{Hom}_{S_{dg}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}},R_{\mathbf{q}})$
is a trivial in $D(k)$ does not hold. Thus $\mathrm{Hom}_{S_{dg}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}},R_{\mathbf{q}})$ is a nontrivial in $D(k)$. From the isomorphism
$\mathrm{Hom}_{D^{b}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}},R_{\mathbf{q}})\cong\left(\mathrm{Hom}_{D^{b}(R^{e})}(R,R)\right)_{\mathbf{q}},$
as algebras in the above argument, we know that $\mathbf{q}$ is contained in $\mathrm{Supp}(S_{dg}(R))$, which suggests that
$V(I)\subseteq\mathrm{Supp}(S_{dg}(R)).$
Thus we obtain $V(I)=\mathrm{Supp}(S_{dg}(R))$, that is, the support scheme of $\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)$ in $D(k)$ is $V(I)$. We thus completed the proof. ∎
Via the algebra morphism $\iota_{I}$, we view $\mathrm{HH}^{0}_{sg}(R)$ as an $R/I$-module.
Lemma 3.9.
With the above setting, $\mathrm{HH}^{0}_{sg}(R)$ is a finitely generated $R/I$-module.
Proof.
First, we view $\mathrm{HH}^{0}_{sg}(R)$ as an $R$-module by the algebra morphism $\pi_{R}:R\rightarrow\mathrm{HH}^{0}_{sg}(R)$. Now, by Proposition 3.6, we have $\mathrm{HH}_{sg}^{0}(R)=\mathrm{H}^{0}\big{(}\mathrm{Cone}(E^{\ast}\rightarrow F^{*})\big{)}$.
Second, since $R$ is a Gorenstein Noetherian algebra, $R^{e}$ is also a Gorenstein Noetherian algebra. Hence, $R^{e}$ admits a bounded finitely generated injective $R^{e}$-module resolution, denoted by $J^{*}$. There is a quasi-isomorphism of $R^{e}$-modules complexes:
$\mathbb{R}\mathrm{Hom}_{R}(\mathrm{Bar}(R),J^{*})\rightarrow\mathbb{R}\mathrm{Hom}_{R}(\mathrm{Bar}(R),R^{e})$
induced by the injective resolution $R^{e}\hookrightarrow J^{*}$ and another quasi-isomorphism of $R^{e}$-modules complexes:
$\mathbb{R}\mathrm{Hom}_{R}(R,J^{*})\rightarrow\mathbb{R}\mathrm{Hom}_{R}(\mathrm{Bar}(R),J^{*})$
induced by Bar resolution $\mathrm{Bar}(R)\twoheadrightarrow R$. Since $\mathbb{R}\mathrm{Hom}_{R}(R,J^{*})$ is a bounded finitely generated $R^{e}$-module complex, it admits a finitely generated projective $R^{e}$-module resolution, denoted by $P^{\ast}_{J^{*}}$. It suggests that there are quasi-isomorphisms
$\mathrm{Bar}(R)\otimes_{R^{e}}P^{\ast}_{J^{*}}\xrightarrow{\sim}\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R}(R,J^{*})\xrightarrow{\sim}\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R}(\mathrm{Bar}(R),J^{*})$
$\xrightarrow{\sim}\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R}(\mathrm{Bar}(R),R^{e})\xrightarrow{\sim}E^{*}.$
Moreover, there is a quasi-isomorphism
$\mathrm{Bar}(R)\otimes_{R^{e}}P^{\ast}_{J^{*}}\xrightarrow{\sim}R\otimes_{R^{e}}P^{\ast}_{J^{*}}.$
It follows that
$R\otimes_{R^{e}}P^{\ast}_{J^{*}}\cong E^{*}$
in $D(k)$. Meanwhile, we know that $P^{j}_{J^{*}}$ is a finitely generated projective $R^{e}$-module for any $j$. Thus, $R\otimes_{R^{e}}P^{j}_{J^{*}}$ is a finitely generated $R$-module for any $j$. Since $R\otimes_{R^{e}}P^{\ast}_{J^{*}}$ is a finitely generated $R$-module complex, the cohomologies of $E^{\ast}$ are all finitely generated $R$-modules.
From the distinguished triangle
$F^{\ast}\rightarrow\mathrm{Cone}(E^{\ast}\rightarrow F^{*})\rightarrow E^{\ast}[1],$
we have the long exact sequence
$\cdots\rightarrow\mathrm{H}^{0}(E^{\ast})\rightarrow\mathrm{H}^{0}(F^{\ast})\xrightarrow{\pi_{R}}\mathrm{H}^{0}\big{(}\mathrm{Cone}(E^{\ast}\rightarrow F^{*})\big{)}\rightarrow\mathrm{H}^{1}(E^{\ast})\rightarrow\cdots,$
where $\mathrm{H}^{0}(F^{\ast})\cong\mathrm{HH}^{0}(R)\cong R$ and $\mathrm{H}^{0}\big{(}\mathrm{Cone}(E^{\ast}\rightarrow F^{*})\big{)}\cong\mathrm{HH}_{sg}^{0}(R)$. Since both $\mathrm{H}^{0}(F^{\ast})$ and $\mathrm{H}^{1}(E^{\ast})$ are finitely generated $R$-modules, we get that $\mathrm{HH}_{sg}^{0}(R)$ is also a finitely generated $R$-module from the long exact sequence.
Since $\pi_{R}$ factors through algebra morphism $\iota_{I}:R/I\hookrightarrow\mathrm{HH}_{sg}^{0}(R)$ (see Proposition 3.8), we obtain that $\mathrm{HH}_{sg}^{0}(R)$ is a finitely generated $R/I$-module. We thus completed the proof. ∎
4. Proof of the main result
In this section, we prove Theorem 1.1. First, we prove the following.
Lemma 4.1.
Let $R$ be a commutative Gorenstein Noetherian ring. Then $D_{sg}(R)$ has a unique DG enhancement.
Proof.
In [22], Rouquier showed that, for any separated scheme $Z$ of finite type over $k$, $D^{b}(Z)$ admits a strong generator. Meanwhile, we know that any strong generator must be a classical generator. Thus suppose $Z$ is an affine scheme $\mathrm{Spec}(R)$, then $D^{b}(R)$ admits a classical generator, denoted by $E$.
Next, the natural localization functor $\pi:D^{b}(R)\rightarrow D^{b}(R)\big{/}\mathrm{Perf}(R)=D_{sg}(R)$ maps $E$ to a classical generator $\pi(E)$ of $D_{sg}(R)$. Meanwhile, since $R$ is Gorenstein, we have
$D_{sg}(R)\simeq\underline{\mathrm{CM}}(R)$
as triangulated categories. Then $D_{sg}(R)$ is an algebraic triangulated category. Thus we have the following triangle equivalence:
$D_{sg}(R)\xrightarrow{\sim}\mathrm{Perf}(T),$
where $T:=\mathbb{R}\mathrm{Hom}_{D_{sg}(R)}(E,E)$ (see [14]).
Note that $T$ is a differential graded algebra. By applying [18, Theorem 2.8], we know that $D(T)^{c}=\mathrm{Perf}(T)$ has a unique DG enhancement, where $(D(T))^{c}$ is the full subcategory, consisting of compact objects, of $D(T)$. It implies that $D_{sg}(R)$ has a unique DG enhancement which is given by the Drinfeld quotient. We completed the proof. ∎
By the above lemma, we know that $S_{dg}(R)$ is the unique DG enhancement of $D_{sg}(R)$.
Lemma 4.2.
Let $R_{1}$ and $R_{2}$ be two commutative Gorenstein Noetherian rings. Suppose that there is a singular equivalence
$F:D_{sg}(R_{1})\rightarrow D_{sg}(R_{2}).$
Then there is a quasi-equivalence
$\tilde{F}:S_{dg}(R_{1})\rightarrow D_{sg}(R_{2}),$
where $\tilde{F}$ is a lift of $F$, i.e., $F=\mathrm{H}(\tilde{F})$.
Proof.
By Lemma 4.1, $\mathrm{Perf}(T)$ has a unique DG enhancement. Thus, by Definition 2.8, there is a quasi-equivalence
$\tilde{F}:S_{dg}(R_{1})\rightarrow S_{dg}(R_{2})$
such that it is a lift of $F$, i.e., $F=\mathrm{H}(\tilde{F})$ (see [2, 1.15]). QED. ∎
Now we are ready to show the following.
Proof of Theorem 1.1.
Retain the settings in Proposition 3.8 and Lemma 3.9. Since the algebra morphism $\iota_{I}$ gives the $R/I$-module structure on $\mathrm{HH}_{sg}^{0}(R)$, the image of $\iota_{I}$ is contained in the central $Z\big{(}\mathrm{HH}_{sg}^{0}(R)\big{)}$ of $\mathrm{HH}_{sg}^{0}(R)$. Now, let
$(\iota_{I})_{\natural}:R/I\rightarrow Z\big{(}\mathrm{HH}_{sg}^{0}(R)\big{)}.$
be the induced map. By Lemma 3.9, it is clear that $(\iota_{I})_{\natural}$ is a finite morphism of Noetherian rings, i.e., $Z\big{(}\mathrm{HH}_{sg}^{0}(R)\big{)}$ is a finitely generated $R/I$-module. Hence, $Z\big{(}\mathrm{HH}_{sg}^{0}(R)\big{)}$ is also a Noetherian ring. By Proposition 3.8, we know that $(\iota_{I})_{\natural}$ is also an injection. It follows that there exists a surjective morphism of schemes:
$(\tilde{\iota}_{I})_{\natural}:\mathrm{Spec}\left(Z\big{(}\mathrm{HH}_{sg}^{0}(R)\big{)}\right)\twoheadrightarrow\mathrm{Spec}(R/I)$
given by $(\iota_{I})_{\natural}$. Moreover, $(\tilde{\iota}_{I})_{\natural}$ is also a finite morphism between schemes since $(\iota_{I})_{\natural}$ is a finite morphism of Noetherian rings. We thus completed the proof. ∎
From the above theorem, we obtain the following.
Corollary 4.3.
Let $R_{1}$ and $R_{2}$ be two commutative Gorenstein Noetherian rings. Suppose that there is a singular equivalence
$D_{sg}(R_{1})\rightarrow D_{sg}(R_{2}).$
Then the two singular loci in the affine schemes of these two rings have the same dimension.
Proof.
Again retain the settings in Proposition 3.8 and Lemma 3.9. Since $(\tilde{\iota}_{I})_{\natural}$ is a finite morphism between schemes, $(\tilde{\iota}_{I})_{\natural}$ has finite fibers, i.e., any one of its fibers has finite points. It implies that
$\mathrm{dim}\big{(}\mathrm{Spec}(R/I)\big{)}=\mathrm{dim}\left(\mathrm{Spec}\left(Z\big{(}\mathrm{HH}_{sg}^{0}(R)\big{)}\right)\right)$
since $(\tilde{\iota}_{I})_{\natural}$ is surjective. Thus the dimension of singular locous of $\mathrm{Spec}(R)$ is equal to the Krull dimension of $Z\big{(}\mathrm{HH}_{sg}^{0}(R)\big{)}$.
By Lemma 4.2, we obtain that there is a quasi-equivalence
$\tilde{F}:S_{dg}(R_{1})\xrightarrow{(-)\otimes^{\mathbb{L}}_{R_{1}}M}S_{dg}(R_{2})$
induced by the triangulated functor $F$. Thus by Theorem 3.3, their Hochschild cohomologies are isomorphic. It implies that
$Z\big{(}\mathrm{HH}_{sg}^{0}(R_{1})\big{)}\cong Z\big{(}\mathrm{HH}_{sg}^{0}(R_{2})\big{)}$
as algebras. Therefore, the two singular loci in these two schemes of $R_{1}$ and $R_{2}$ respectively, have the same dimension. ∎
Acknowledgements.
We would like to thank Xiaojun Chen and Huijun Fan for their encouragement, support and suggestions.